# FedFitTech: A Baseline in Federated Learning for Fitness Tracking

### v1

@@ Abstract.
Rapid evolution of sensors and resource-efficient machine learning models have spurred the widespread adoption of wearable fitness tracking devices. Equipped with inertial sensors, such devices can continuously capture physical movements for fitness technology (FitTech), enabling applications from sports optimization to preventive healthcare. Traditional centralized learning approaches to detect fitness activities struggle with data privacy concerns, regulatory constraints, and communication inefficiencies. In contrast, Federated Learning enables a decentralized model training by communicating model updates rather than potentially private wearable sensor data. Applying Federated Learning to FitTech presents unique challenges, such as data imbalance, lack of labelled data, heterogeneous user activity patterns, and trade-offs between personalization and generalization. To simplify research on FitTech in Federated Learning, we present the FedFitTech baseline, under the Flower framework, which is publicly available and widely used by both industry and academic researchers. Additionally, to illustrate its usage, this paper presents a case study that implements a system based on the FedFitTech baseline, incorporating a client-side early stopping strategy and comparing the results. For instance, this system allows wearable devices to optimize the trade-off between capturing common fitness activity patterns and preserving individuals’ nuances, thereby enhancing both the scalability and efficiency of privacy-aware fitness tracking applications. Results show that this reduces overall redundant communications by 13%, while maintaining the overall recognition performance at a negligible recognition cost by 1%. Thus, FedFitTech baseline creates a foundation for a wide range of new research and development opportunities in FitTech, and it is available as open-source at: https://github.com/adap/flower/tree/main/baselines/fedfittech
@@ 1. Introduction
With recent advancements in sensor technologies and resource-efficient machine learning models, wearable fitness trackers in various form factors such as wrist watches, glasses, earbuds, and rings have gained widespread use. These devices generate vast amounts of valuable sensor data for Human Activity Recognition (HAR) applications, in which body movements can be analyzed in detail with minimal intrusion. For example, inertial sensors can continuously monitor motion and gestures at specific body locations, thus providing a detailed representation of the user’s motion patterns. This makes such devices particularly useful for an extensive range of applications in fitness technology (FitTech), including medical assistance and the optimization of complex work processes (Bulling et al., 2014).
@@
FitTech is a crossover term that combines fitness and technology, encompassing innovations that enhance fitness, wellness, and health experiences. Using wearable devices, mobile apps, and virtual personal training platforms, FitTech can monitor physical activity, provide real-time feedback to user, and can tailor exercise programs. Through machine learning, FitTech promises to transform the fitness industry by making workouts more personalized, efficient, and data-driven, helping users achieve better training results.
@@
Current research tends to focus largely on centralized learning of fitness activities, which can be used to develop potent machine learning models by being able to process large datasets and leverage substantial computational power. However, centralized learning requires sensor data to be shared and stored in a central location, which can deter data sharing due to regulatory restrictions (e.g., the European General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017) and the California Consumer Privacy Act (CCPA) (de la Torre, 2018)), making users reluctant to grant access for training. Moreover, centralized learning struggles to meet the needs for scalability and communication efficiency.
@@
To address these challenges, Federated Learning (McMahan et al., 2017) has emerged as an alternative to centralized learning, offering a model communication mechanism instead of raw data sharing. In Federated Learning, a global model is shared to devices (also known as clients, nodes, or workers) for local training using local data. These tuned local models are then sent back to the server, which aggregates them to update the global model. The updated global model is redistributed to the devices for further training rounds. By the end of the process, all devices obtain a global model trained on all local data, while data privacy is maintained as no sensitive raw data is transmitted to external sources (see Fig. 1). Additionally, Federated Learning offers benefits such as improved data availability, scalability, fault recovery, and communication efficiency (Liu et al., 2021).
@@
While Federated Learning is widely recognized as a viable approach for training machine learning models on decentralized devices, its potential impact on FitTech remains underexplored. Wearables such as smartwatches typically share similar sensors, in particular MEMS inertial sensors, as well as processing and wireless networking components, making FitTech applications relatively homogeneous compared to other environments. This similarity results in consistent computation and communication capabilities across smartwatches, allowing to train the same global model with synchronized communication.
@@
Additionally, McMahan et al (McMahan et al., 2017) identified communication as the primary challenge in Federated Learning systems. However, unlike other domains, fitness tracking does not require daily communication. Since people do not exercise frequently throughout the day, data is not generated continuously, and wearable devices accumulate small datasets. This allows for less frequent and more convenient communication times, such as during charging or when connected to unmetered Wi-Fi. Moreover, modern processors make computing tiny machine learning models more efficient, reducing costs. Furthermore, Ek et al. (Ek et al., 2022) compared various Federated Learning algorithms and found that Federated Learning can achieve comparable or even superior performance to Centralized Learning techniques in the HAR domain. Thus, the FitTech environment is naturally well-suited to Federated Learning.
@@
On the other hand, each person exhibits unique movement patterns, and the performing of the same exercise can vary depending on individual performance. The position of sensors on a person can affect the motion patterns that are recorded, even when repeating the same actions (Arikumar et al., 2022). Factors such as gender, age, weight, height, and exercise frequency further contribute to these unique features (Li et al., 2021). For instance, one person might jog once a week, while another person might jog every other day, leading eventually to data imbalances. Literature shows that Federated Learning underperforms when device data are imbalanced, as excessive external patterns can reduce performance on local data due to the personal nature of movements (Cheng et al., 2023). Moreover, annotated local data are often limited, requiring other people’s patterns to accurately detect a person’s exercise. Given the advantages of Federated Learning in FitTech, we thus present in this paper FedFitTech, a Federated Learning baseline designed for research in the FitTech domain.
@@
Furthermore, we give an example using the FedFitTech baseline by designing a client-side early-stopping implementation and analyze its benefits. In FitTech, due to varied user motion patterns, the global model should not override individual variations. Thus, early exit from FL training may be necessary when the global model underperforms on local datasets. The attractiveness for implementing client-side early stopping in FedFitTech is that in the standard Federated Learning setting (McMahan et al., 2017), all devices participate in the training of the global model for a fixed number of global rounds. This can lead to unnecessary resource consumption for devices that do not need to attend all training rounds. In addition, the FitTech domain requires incorporating only common patterns from others without losing the influence of personal data in the global model. Although Federated Learning provides a private way to benefit from external patterns, users need the global model to be tailored to local data. Therefore, it is essential to design a Federated Learning system that balances personalization and generalization in FitTech. Thus, our case study of FedFitTech baseline involves an early stopping method to mitigate these challenges.
Our contributions are as threefold:
@@
  • Use Case Experiment of Early Stopping: We present an example usage of this baseline, by allowing client-side early stopping to reduce energy consumption and improving efficiency.
@@ 2. Related Work
Several studies have investigated various implementations of Federated Learning for HAR, addressing different challenges through novel strategies. FL-PMI (Arikumar et al., 2022) proposes a system leveraging unlabeled data by applying a cleaning process with auto-labeling propagation to support smart healthcare applications. Meta‐HAR (Li et al., 2021) employs a model‐agnostic meta-learning approach that extracts adaptive signal embeddings from sensor data and integrates a task‐specific classification network via fine-tuning. PMF (Feng et al., 2020) adopts a secure data handling approach by combining differential privacy with secure multi-party computation and introduces an adaptive weight adjustment mechanism to ensure fair contributions to a global model. Also, the work by (Xiao et al., 2021) presents a perceptive extraction network system that enhances data representation while employing encryption and decryption to mitigate potential data leakage risks.
@@
To address challenges of inadequate data and secure data sharing, researchers in (Zhou et al., 2022a) propose a two-dimensional framework that employs both horizontal and vertical FL paradigms, while another study (Zhou et al., 2022b) utilises specialized training algorithms aimed at increasing data diversity and fostering inclusivity for devices with inferior networking conditions.
@@
Focusing on model training improvements, several works explore client grouping and dynamic layer sharing. PS-PFL (Chai et al., 2024) introduces a personalized FL model based on profile similarity that aggregates local models through weighted computations. An attention-based clustering mechanism in (Bu et al., 2024) calculates an optimal weighted model combination to encourage similar clients to group together without relying on a single global model. ClusterFL (Ouyang et al., 2021) leverages a cluster indicator matrix and iterative optimization techniques to update client model weights, facilitating cluster-wise straggler dropout and correlation-based client selection. ProtoHAR (Cheng et al., 2023) integrates a prototype-guided technique that separates representation learning from categorization, thereby achieving uniform activity feature representation and improved convergence. Also, dynamic layer-sharing are investigated by FedCLAR (Presotto et al., 2022), which selectively shares portions of model weights based on a hierarchical clustering, and by FedDL (Tu et al., 2021), which merges lower layers of models to capture user-specific similarities while enhancing communication efficiency.
@@
Additionally, to enhance privacy, the researcher of PDP-FL (Personalized Differential Privacy-based Federated Learning) (Shao, 2024) specifically investigated fitness tracking application using two stages. First, it allows users to customize their privacy levels, with corresponding noise added to their local model prior to sharing with a server. The second stage involves the server adding further noise to ensure global privacy protection.
@@
Weaknesses noted across these studies include that FL-PMI (Arikumar et al., 2022) may require additional computational resources, which could impact battery life, and does not explicitly address sensor inaccuracies. Meta‐HAR (Li et al., 2021) may incur higher computational and communication overhead, posing challenges for very large-scale deployments. While promising, PMF (Feng et al., 2020) requires further empirical validation of its performance under large-scale data scenarios and the integration of scalable privacy-preserving measures. The two-dimensional framework by (Zhou et al., 2022a) could potentially expose finer-grained activity details under adversarial conditions, and the approach in (Zhou et al., 2022b) may face challenges with dynamic client participation and scaling to real-world deployments. Additionally, client grouping strategies such as PS-PFL (Chai et al., 2024) incur extra computational overhead on resource-constrained devices, and ClusterFL (Ouyang et al., 2021) might expose user-cluster associations under adversarial conditions while necessitating an additional optimization step. Lastly, ProtoHAR (Cheng et al., 2023) assumes a stationary data distribution, which might limit its adaptability in continuously evolving environments, FedCLAR (Presotto et al., 2022) assumes the availability of ample labeled data—a condition that may not hold universally—and FedDL (Tu et al., 2021) requires further integration of robust privacy measures along with additional validation of its effectiveness beyond HAR. Besides, while PDP-FL (Shao, 2024) enhances privacy by allowing varied user privacy preferences, this diversity can impact model performance and effectiveness. In other words, differing noise levels, corresponding to individual privacy choices, influence global model aggregation. To add more, PDP-FL simulate fitness tracking environment using image data, which does not fully capture the complexities of the domain. Fitness data often has temporal dependencies, varying sampling rates, and different types of noise characteristics compared to image pixels.
@@
While other studies addressed HAR applications, this study especially focuses on the FitTech domain using FL, similar to PDP-FL. However, our FedFitTech baseline utilizes a real fitness tracking dataset and a tiny model, reflecting original FitTech characteristics and providing further exploration of FL potantial in this area.
@@
About designing case study of FedFitTech, the implementation of early stopping in the FL setting has already been explored in FLrce (Niu et al., 2024) and FLASH (Panchal et al., 2023). FLrce employs a resource-efficient early-stopping strategy tailored for FL. FLASH leverages concept drift detection through client-side early stopping and server-side adaptive optimization. However, their complex mechanisms increase computational overhead and risk premature termination. Also, these methods are primarily designed for image recognition in non-FitTech domains. Our case study especially focuses on the FitTech domain and employs an early-stopping approach based on the stability of the validation F1-score over a sliding window (Prechelt, 2002). While this method is simple to implement, it ensures that the global model retains valuable patterns learned from all clients, including those with less significant contributions, which is an essential factor in FL for FitTech. Moreover, case study emphasizes communication efficiency, and balanced generalization. Despite a neglible F1-score drop, our case study allows clients to benefit from extarnal patterns and terminate training when global weights diverge from local patterns, preventing redundant computation and communication.
@@ 3.1. Model
Over the past decade, Deep Learning (DL) methods have become the dominant approach in inertial-based HAR, consistently outperforming traditional Machine Learning algorithms (Ordóñez and Roggen, 2016; Guan and Plötz, 2017). For this reason, we used TinyHAR (Zhou et al., 2022c) which is a lightweight DL model specifically designed for HAR. TinyHAR leverages a combination of convolutional layers, a transformer block featuring self-attention and fully connected layers, and a Long Short-Term Memory (LSTM) to extract and model features from sensor data efficiently. Initially, convolutional layers are applied to each sensor channel to capture local patterns. This is followed by a transformer block that facilitates effective cross-channel feature interaction. Next, a fully connected layer fuses the extracted features, which are then processed by an LSTM to capture long-term temporal dependencies. Finally, a self-attention layer dynamically recalibrates the importance of features across time. This architecture achieves competitive performance with a significantly reduced parameter count, making it highly suitable for deployment on resource-constrained devices.

### v2

@@ Abstract.
The rapid evolution of sensors and resource-efficient machine learning models has spurred the widespread adoption of wearable fitness tracking devices. Equipped with inertial sensors, such devices can continuously capture physical movements for fitness technology (FitTech), enabling applications from sports optimization to preventive healthcare. Traditional Centralized Learning approaches to detect fitness activities struggle with data privacy concerns, regulatory restrictions, and communication inefficiencies. In contrast, Federated Learning enables a decentralized model training by communicating model updates rather than potentially private wearable sensor data. Applying Federated Learning to FitTech presents unique challenges, such as data imbalance, lack of labeled data, heterogeneous user activities, and trade-offs between personalization and generalization. To simplify research on FitTech in Federated Learning, we present the FedFitTech baseline, under the Flower framework, which is publicly available and widely used by both industry and academic researchers. Additionally, to illustrate its usage, this paper presents a case study that implements a system based on the FedFitTech baseline, incorporating a client-side early stopping strategy and comparing the results. For instance, this system allows wearable devices to optimize the trade-off between capturing common fitness activities and preserving individuals’ nuances, thereby enhancing both the scalability and efficiency of privacy-aware fitness tracking applications. The results show that this reduces the overall redundant communications by 13%, while maintaining the overall recognition performance at a negligible recognition cost by 1%. Thus, the FedFitTech baseline creates a foundation for a wide range of new research and development opportunities in FitTech, and it is available as open source at: https://github.com/shreyaskorde16/FedFitTech
@@ 1. Introduction
With recent advances in sensor technologies and resource-efficient machine learning models, wearable fitness trackers in various forms, such as wrist watches, glasses, earbuds, and rings, have gained widespread use. These devices generate a vast amount of valuable sensor data for Human Activity Recognition (HAR) applications, in which body movements can be analyzed in detail with minimal intrusion. For example, inertial sensors can continuously monitor motion and gestures at specific body locations, thus providing a detailed representation of the user’s motion patterns. This makes such devices particularly useful for an extensive range of applications in fitness technology (FitTech), including medical assistance and the optimization of complex work processes (Bulling et al., 2014).
@@
FitTech is a crossover term that combines fitness and technology, encompassing innovations that enhance fitness, wellness, and health experiences. Using wearable devices, mobile apps, and virtual personal training platforms, FitTech can monitor physical activity, provide real-time feedback to users, and tailor exercise programs to their needs. Through machine learning, FitTech promises to transform the fitness industry by making workouts more personalized, efficient, and data-driven, helping users achieve better training results.
@@
Current research tends to focus largely on Centralized Learning of fitness activities, which can be used to develop powerful machine learning models by being able to process large datasets and leverage substantial computational power. However, Centralized Learning requires sensor data to be shared and stored in a central location, which can deter data sharing due to regulatory restrictions (e.g., the European General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017) and the California Consumer Privacy Act (CCPA) (de la Torre, 2018)), making users reluctant to grant access for training. Moreover, Centralized Learning struggles to meet the needs for scalability and communication efficiency.
@@
To address these challenges, Federated Learning (McMahan et al., 2017) has emerged as an alternative to Centralized Learning, offering a model communication mechanism instead of raw data sharing. In Federated Learning, a global model is shared with devices (also known as clients, nodes, or workers) for local training using local data. These tuned local models are then sent back to the server, which aggregates them to update the global model. The updated global model is redistributed to the devices for further training rounds. At the end of the process, all devices obtain a global model trained on all local data, while data privacy is maintained, as sensitive raw data are not transmitted to external sources (see Fig. 1). Additionally, Federated Learning offers benefits such as improved data availability, scalability, fault recovery, and communication efficiency (Liu et al., 2021).
@@
Although Federated Learning is widely recognized as a viable approach for training machine learning models on decentralized devices, its potential impact on FitTech remains underexplored. Wearables such as smartwatches typically share similar sensors, particularly MEMS inertial sensors, as well as processing and wireless networking components, making FitTech applications relatively homogeneous compared to other environments. This similarity results in consistent computation and communication capabilities across smartwatches, enabling training of the same global model with synchronized communication.
@@
Additionally, researchers of (McMahan et al., 2017) identified communication as the primary challenge in Federated Learning systems. However, unlike other domains, fitness tracking does not require daily communication. Since people do not exercise frequently throughout the day, data is not generated continuously, and wearable devices accumulate small datasets. This allows for less frequent and more convenient communication times, such as during charging or when connected to unmetered Wi-Fi. Moreover, modern processors make the computation of tiny machine learning models more efficient, reducing costs. Furthermore, authors of (Ek et al., 2022) compared various Federated Learning algorithms and found that Federated Learning can achieve comparable or even superior performance to Centralized Learning techniques in the HAR domain. Thus, the FitTech environment is naturally well suited to Federated Learning.
@@
On the other hand, each person exhibits unique movement patterns and the performing of the same exercise can vary depending on individual performance. The position of the sensors on a person can affect the recorded motion patterns, even when repeating the same actions (Arikumar et al., 2022). Factors such as gender, age, weight, height, and exercise frequency further contribute to these unique characteristics (Li et al., 2021). For instance, one person might jog once a week, while another person might jog every other day, eventually leading to data imbalances. The literature shows that Federated Learning underperforms when device data is imbalanced, as excessive external patterns can reduce performance on local data due to the personal nature of movements (Cheng et al., 2023). Moreover, annotated local data is often limited, requiring patterns from other people to accurately detect a person’s exercise. Given the advantages of Federated Learning in FitTech, we thus present in this paper FedFitTech, a Federated Learning baseline designed for research in the FitTech domain.
@@
Furthermore, we give an example using the FedFitTech baseline by designing a client-side early stopping implementation and analyzing its benefits. In FitTech, due to the varied motion patterns of the user, the global model should not override individual variations. Thus, early stopping may be helpful when the global model underperforms on local datasets. The attractiveness of implementing client-side early stopping in FedFitTech is that, in the standard Federated Learning setting (McMahan et al., 2017), all devices participate in the training of the global model for a fixed number of communication rounds. This can lead to unnecessary resource consumption for devices that do not need to attend all training rounds. In addition, the FitTech domain requires incorporating only common patterns from others without losing the influence of personal data in the global model. Although Federated Learning provides a private way to benefit from external patterns, users need the global model to be tailored to local data. Therefore, it is essential to design a Federated Learning system that balances personalization and generalization in FitTech. Thus, our case study of the FedFitTech baseline involves an early stopping method to mitigate these challenges.
Our contributions are threefold:
@@
  • Use Case Experiment of Early Stopping: We present an example usage of this baseline, by allowing client-side early stopping to reduce energy consumption and thus improving the overall efficiency of the system.
@@ 2. Related Work
Several studies have investigated various implementations of Federated Learning for HAR, addressing different challenges through novel strategies. FL-PMI (Arikumar et al., 2022) proposes a system leveraging unlabeled data by applying a cleaning process with auto-labeling propagation to support smart healthcare applications. Meta‐HAR (Li et al., 2021) employs a model‐agnostic meta-learning approach that extracts adaptive signal embeddings from sensor data and integrates a task‐specific classification network through fine-tuning. PMF (Feng et al., 2020) adopts a secure data handling approach by combining differential privacy with secure multi-party computation and introduces an adaptive weight adjustment mechanism to ensure fair contributions to a global model. Also, the work by (Xiao et al., 2021) presents a perceptive extraction network system that enhances data representation while employing encryption and decryption to mitigate potential data leakage risks.
@@
To address the challenges of inadequate data and secure data sharing, researchers in (Zhou et al., 2022a) propose a two-dimensional framework that employs both horizontal and vertical FL paradigms, while another study (Zhou et al., 2022b) utilizes specialized training algorithms aimed at increasing data diversity and fostering inclusivity for devices with inferior network conditions.
@@
Focusing on model training improvements, several works explore client grouping and dynamic layer sharing. PS-PFL (Chai et al., 2024) introduces a personalized FL model based on profile similarity that aggregates local models through weighted computations. An attention-based clustering mechanism in (Bu et al., 2024) calculates an optimal weighted model combination to encourage similar clients to group together without relying on a single global model. ClusterFL (Ouyang et al., 2021) leverages a cluster indicator matrix and iterative optimization techniques to update client model weights, facilitating cluster-wise straggler dropout and correlation-based client selection. ProtoHAR (Cheng et al., 2023) integrates a prototype-guided technique that separates representation learning from categorization, thereby achieving uniform activity feature representation and improved convergence. Also, dynamic layer-sharing is investigated by FedCLAR (Presotto et al., 2022), which selectively shares portions of model weights based on a hierarchical clustering, and by FedDL (Tu et al., 2021), which merges lower layers of models to capture user-specific similarities while enhancing communication efficiency.
@@
Furthermore, to enhance privacy, the researcher of PDP-FL (Personalized Differential Privacy-based Federated Learning) (Shao, 2024) specifically investigated the fitness tracking application using two stages. First, it allows users to customize their privacy levels, with corresponding noise added to their local model prior to sharing with a server. The second stage involves the server adding further noise to ensure global privacy protection.
@@
The weaknesses noted across these studies include that FL-PMI (Arikumar et al., 2022) may require additional computational resources, which could impact battery life, and does not explicitly address sensor inaccuracies. Meta‐HAR (Li et al., 2021) may incur higher computational and communication overhead, posing challenges for very large-scale deployments. While promising, PMF (Feng et al., 2020) requires further empirical validation of its performance under large-scale data scenarios and the integration of scalable privacy-preserving measures. The two-dimensional framework by (Zhou et al., 2022a) could potentially expose finer-grained activity details under adversarial conditions, and the approach in (Zhou et al., 2022b) may face challenges with dynamic client participation and scaling to real-world deployments. Additionally, client grouping strategies such as PS-PFL (Chai et al., 2024) incur extra computational overhead on resource-constrained devices, and ClusterFL (Ouyang et al., 2021) might expose user-cluster associations under adversarial conditions while necessitating an additional optimization step. Lastly, ProtoHAR (Cheng et al., 2023) assumes a stationary data distribution, which might limit its adaptability in continuously evolving environments, FedCLAR (Presotto et al., 2022) assumes the availability of ample labeled data — a condition that may not hold universally — and FedDL (Tu et al., 2021) requires further integration of robust privacy measures along with additional validation of its effectiveness beyond HAR. Besides, while PDP-FL (Shao, 2024) enhances privacy by allowing varied user privacy preferences, this diversity can impact model performance and effectiveness. In other words, differing noise levels, corresponding to individual privacy choices, influence global model aggregation. To add more, PDP-FL simulates the fitness tracking environment using image data, which does not fully capture the complexities of the domain. Fitness data often have temporal dependencies, varying sampling rates, and different types of noise characteristics compared to image pixels.
@@
While other studies addressed HAR applications, this study especially focuses on the FitTech domain using FL, similar to PDP-FL. However, our FedFitTech baseline utilizes a real fitness tracking dataset and a tiny model, reflecting original FitTech characteristics and providing further exploration of FL potential in this area.
@@
About designing a case study of FedFitTech, the implementation of early stopping in the FL setting has already been explored in FLrce (Niu et al., 2024) and FLASH (Panchal et al., 2023). FLrce employs a resource-efficient early stopping strategy tailored for FL. FLASH leverages concept drift detection through client-side early stopping and server-side adaptive optimization. However, their complex mechanisms increase computational overhead and risk premature termination. Also, these methods are primarily designed for image recognition in non-FitTech domains. Our case study especially focuses on the FitTech domain and employs an early stopping approach based on the stability of the validation F1-score over a sliding window (Prechelt, 2002). While this method is simple to implement, it ensures that the global model retains valuable patterns learned from all clients, including those with less significant contributions, which is an essential factor in FL for FitTech. Moreover, the case study emphasizes communication efficiency and balanced generalization. Despite a negligible F1-score drop, our case study allows clients to benefit from external patterns and terminate training when global weights diverge from local patterns, preventing redundant computation and communication.
@@ 3.1. Model
Over the past decade, Deep Learning (DL) methods have become the dominant approach in inertial-based HAR, consistently outperforming traditional Machine Learning algorithms (Ordóñez and Roggen, 2016; Guan and Plötz, 2017). For this reason, we used TinyHAR (Zhou et al., 2022c), which is a lightweight DL model specifically designed for HAR. TinyHAR leverages a combination of convolutional layers, a transformer block featuring self-attention and fully connected layers, and a Long Short-Term Memory (LSTM) to extract and model features from sensor data efficiently. Initially, convolutional layers are applied to each sensor channel to capture local patterns. This is followed by a transformer block that facilitates effective cross-channel feature interaction. Next, a fully connected layer fuses the extracted features, which are then processed by an LSTM to capture long-term temporal dependencies. Finally, a self-attention layer dynamically recalibrates the importance of features across time. This architecture achieves competitive performance with a significantly reduced parameter count, making it highly suitable for deployment on resource-constrained devices.